### v1

# Coherence-Driven Multimodal Safety Dialogue with Active Learning for Embodied Agents

Abstract.
When assisting people in daily tasks, robots need to accurately interpret visual cues and respond effectively in diverse safety-critical situations, such as sharp objects on the floor. In this context, we present M-CoDAL, a multimodal-dialogue system specifically designed for embodied agents to better understand and communicate in safety-critical situations. The system leverages discourse coherence relations to enhance its contextual understanding and communication abilities. To train this system, we introduce a novel clustering-based active learning mechanism that utilizes an external Large Language Model (LLM) to identify informative instances. Our approach is evaluated using a newly created multimodal dataset comprising 1K safety violations extracted from 2K Reddit images. These violations are annotated using a Large Multimodal Model (LMM) and verified by human annotators. Results with this dataset demonstrate that our approach improves resolution of safety situations, user sentiment, as well as safety of the conversation. Next, we deploy our dialogue system on a Hello Robot Stretch robot and conduct a within-subject user study with real-world participants. In the study, participants role-play two safety scenarios with different levels of severity with the robot and receive interventions from our model and a baseline system powered by OpenAI’s ChatGPT. The study results corroborate and extend the findings from automated evaluation, showing that our proposed system is more persuasive and competent in a real-world embodied agent setting.
Key words and phrases:
Embodied Agents, Multimodal Safety, Active Learning, Coherence Theory
1. Introduction
Embodied agents, such as assistive robots designed for household tasks and daily assistance, should perceive and interpret their surroundings and engage in conversations with human users when noticing an unsafe scenario. The advent of powerful multimodal models (202, 2023; Liu et al., 2023) can allow robots to interpret and respond to cues from both visual and conversational inputs in real-world scenarios. However, ensuring user safety in these interactions requires systems that understand and adapt to the conversational context and remain robust across diverse situations. To address this dual challenge, we integrate theories of coherence relations to enhance contextual understanding and employ clustering-based active learning to train a multimodal dialogue system, M-CoDAL, specifically designed for embodied agents to provide safety advice to users.
Figure 1. Our proposed dialogue system first parses safety violation in an image with Penn Discourse Treebank (Prasad et al., 2008) relations, and then generate an appropriate response by choosing an Segmented Discourse Representation Theory relation (Asher and Lascarides, 2005).
M-CoDAL uses discourse coherence relations to enrich its contextual understanding. Theories of coherence relations originate in understanding and analyzing inferential links to support text interpretation and has subsequently been extended to cross-modal settings (Alikhani et al., 2023). Coherence relations can situate an ongoing scene in the arc of a narrative (Cohn, 2013) or enrich interpretation of communicative actions across modalities (Lascarides and Stone, 2009). For instance, the coherence relation Cause can help interpret why paying attention to a safety violation is important in an image captured by a robot given a certain Condition relation. These relations can then facilitate dialogue between the human user and the robot. By default, a robot equipped with pretrained models would not explicitly model such inferential links (Alikhani et al., 2023). We hypothesize that enabling robots to parse coherence relations of safety violation in captured images, and to use these relations to guide their responses, will result in embodied agents that better understand contexts of safety and engage in safer conversations with humans.
To train M-CoDAL for embodied agents with integration of coherence relations, we adopt clustering-based active learning (Settles, 2009). Active learning offers advantage over standard fine-tuning methods by focusing resources toward informative instances. When combined with clustering, active learning has been shown to yield more representative models (Hassan and Alikhani, 2023a). Thus, we hypothesize that applying clustering-based active learning would lead to better coverage of safety scenarios and a safer multimodal dialogue system. Active learning, however, has mostly been addressed in the context of classification tasks and has remained challenging for generative tasks (Perlitz et al., 2023) such as conversational response generation due to difficulty in estimating model uncertainty in a large output space. We make the first inroads in active learning for autonomous agents with conversational capacity by integrating an external LLM that quantifies informativeness of an instance based on a composite score of safety of conversation, resolution, and user sentiment. Further, we distill knowledge from another external LLM to reduce necessary human efforts in active learning. Our work is also one of the first to deploy and evaluate active learning in practice for human-robot interactions.
To train our system and evaluate our approach, we first construct a novel dataset consisting of 1K safety violations obtained from 2K Reddit images. The safety violations are obtained using a Large Multimodal Model (LMM) and then verified, or edited if necessary, by human annotators. Coherence relations and appropriate responses based on the coherence relations are distilled from GPT-4 (202, 2023) for these safety violations and used to train M-CoDAL, based on a smaller LLM, Mistral-7B (Jiang et al., 2023), in an active learning setting. Our automated evaluation demonstrates that integrating coherence relations leads to higher safety scores, further improved by clustering-based active learning. We also observe that the improvements translate to models such as Llama 3 (Dubey et al., 2024) and Qwen (Bai et al., 2023) that were not part of the original active learning loop.
Finally, we deploy our multimodal dialogue system on a Hello Robot Stretch robot and conduct a within-subject user study where real-world participants interact with our robot in household scenarios regarding potential safety violations. 8 participants interact with the robot operating either with GPT-4o or our proposed system, M-CoDAL in a low and high severity fake unsafe scenario. The study results corroborate findings of automated evaluation by showing that participants found our proposed system M-CoDAL to be more persuasive in both conditions. Participants also find M-CoDAL to be more competent in low-serverity situations. Our qualitative analysis of interviews with the participants also reveals that the participants find M-CoDAL to be more attentive to safety situation and can help the user to be more aware of safety situations. Thus, the key contributions of this paper are:
  • A first-of-its-kind publicly available dataset  [^1] https://github.com/sabithsn/multimodal-embodied-safety of multimodal dialogues of safety for embodied agents.
  • A multimodal dialogue system, M-CoDAL that better understands context of safety via coherence relations.
  • Extension of active learning paradigm for conversational embodied agent with integration of an external LLM.
  • Deployment of M-CoDAL with a Hello Robot Stretch robot, accompanied with findings of a real-world user study.
2. Related Work
Safety in Automated Agents
There has been a growing interest in development of multimodal dialogue systems in recent years (Sun et al., 2022a; Lee et al., 2023; Sicilia et al., 2023), especially with the advent of large multimodal models such as GPT-4V (202, 2023) and LLaVA (Liu et al., 2023). Discussions of safety however, has primarily focused on textual domain (Sun et al., 2022b; Weidinger et al., 2022; Dinan et al., 2022; Atwell et al., 2022) and safety of embodied multimodal agents remain relatively unexplored despite its importance. To our knowledge, our work is the first public work to process visual cues of safety in multimodal dialogues with an embodied agent.
Coherence Relations for Contextual Understanding
Coherence relations have been proposed as a possible method for controlling generative models and have been shown to aid tasks such as extractive and abstractive summarization (Cohan et al., 2018; Xu et al., 2020). Coherence relation-aware models have also been shown to generate more coherent texts (Bosselut et al., 2018) within a reinforcement learning setting. Alikhani et al. (2020) extend standard text-based discourse relations to cross-modal scenarios. Our work is the first to leverage coherence relations witin a multimodal dialogue system for safety.
Active Learning with Natural Language
Active learning is a prominent area in machine learning (Settles, 2009), and has gained recent attention for tasks involving natural language (Zhang et al., 2022) such as intent classification, sentence matching, and named entity recognition (Zhang and Zhang, 2019; Bai et al., 2020; Liu et al., 2022). However, active learning has predominantly been focused on classification tasks. Recent works targeting LLMs (Diao et al., 2023; Margatina et al., 2023; Hassan et al., 2024a) also focus on tasks with fixed sets of outputs, leaving active learning for generative tasks largely unexplored (Perlitz et al., 2023; Hassan et al., 2024b). Our work is among the first to apply active learning to generative tasks and multimodal scenarios. Further, we deploy and evaluate active learning in practice, whereas prior work have primarily relied on simulations (Zhang et al., 2022).
Embodied agents with AI models
With advancement in language models and multimodal models, there has been a surge in research integrating these models with embodied agents (Ahn et al., 2022; Driess et al., 2023; Huang et al., 2022b, a). Most of these recent works focus on extracting plans or executable actions by the robot from pre-trained language models. Different from the aforementioned works, ours is the first to develop a multimodal dialogue system for safety in embodied agents and conduct a real-world user study by deploying it with a robot.
3. Multimodal Dialogue Framework
We first outline the setting of our dialogue system M-CoDAL, followed by integration of coherence relations and use of clustering-based active learning for training the dialogue system.
3.1. Dialogue System
The first input to our dialogue system is an image that contains a potential safety violation. These safety violations may occur during household tasks or within a living environment. For the training phase, images from social media are used, and the learner LLM is fine-tuned with four turns of simulated conversation:
Turn #1: A Large Multimodal Model (LMM) processes the image and generates a message describing the safety violation in the image. Turn #2: A user responds to the safety violation issue raised in the first turn. Turn #3: A Large Language Model (LLM) processes the previous two turns and generates a response. Turn #4: The user makes the final response in the conversation.
During the user-study phase, the images are captured by an actual robot in household environment and the conversations are not restricted to a specific number of turns, continuing indefinitely.
3.2. Coherence Relations
In our work, we consider two prominent frameworks for discourse coherence relations (Figure 1). The first is Penn Discourse Treebank (PDTB) (Lin et al., 2014) and the second is Segmented Discourse Representation Theory (SDRT) (Asher and Lascarides, 2005). PDTB coherence relations such as Cause focus on the local relations between adjacent or nearby textual units. SDRT coherence relations such as Background aim to capture semantic and pragmatic discourse structure.
Parsing Safety Violation with PDTB Relations:
Penn Discourse TreeBank (PDTB) is particularly suitable for identifying coherence relations within and across sentences. The safety violation obtained from LMM in Turn #1 is parsed by an external LLM. The LLM is asked to consider PDTB relations that occur more than 1% in intra-sentential scenarios (Liang et al., 2020): Concession, Contrast, Cause, Cause+Belief, Condition, Purpose, Conjunction, Instantiation, Level-of-detail, Manner, Substitution, Asynchronous, Synchronous. The parsed safety violation in Turn #1, along with Turn #2 are passed as context to the LLM to generate Turn #3.
Introducing SDRT Relations in Dialogue:
While generating Turn #3, we let an external LLM decide the appropriate discourse coherence relation to be maintained in the response. Since the coherence relation to be maintained here is at a turn-level, we opt for Segmented Discourse Representation Theory (SDRT) (Asher and Lascarides, 2005) as SDRT has been shown to be effective at turn-level in conversational scenarios (Asher et al., 2016). The external LLM is provided with 16 SDRT relations listed in (Asher et al., 2016): Continuation, Result, Elaboration, Conditional, Contrast, Answer / Question answer pair, Q-elab / Follow-up question, Acknowledgement, Narration, Correction, Explanation, Alternation, Parallel, Commentary, Clarification Q, Background.
Figure 2. Our proposed active learning loop. Informative instances are identified by an external LLM using a composite score. These instances are then distilled using another LLM for discourse relations and responses. These are then used to retrain the learner LLM.
3.3. Active Learning
Preliminaries
We assume there is a pool of unlabeled dataset U but only a subset of labeled data L can be used for training. L is iteratively constructed by querying target output for the most-informative instance. While other active learning scenarios exist (Settles, 2009), we follow the setting of pool-based active learning because of its relevance to our setting, where we can obtain a large number of safety images but obtaining coherence relations and appropriate responses can be challenging.
Active learning for classification tasks typically identifies highly informative instances with measures of uncertainty, such as entropy (Settles, 2009):
[…]
In Eq. 1, i ranges over all possible labels.
Active Learning for Dialogue
Standard measures of informativeness such as entropy, however, cannot be a useful measure in generative setting (Perlitz et al., 2023) such as dialogue systems. This is because, as opposed to classification models, generative models such as LLMs have a massive output space and entropy over such a large space may not indicate informativeness properly. Thus, we propose to replace entropy with a new composite score calculated by an external LLM. To calculate this score, we assume that {$p_{1},p_{2},...p_{k}$} is a set of attributes we expect the generated output to preserve. In our setting, we want the output to be safe and also resolve the safety situation with the user while not upsetting the user. In this case, $p_{1}$ is the safety of the generated response and $p_{2}$ is the resolution score, and $p_{3}$ is the user sentiment. Then, we can define our informative instance as:
[…]
In Eq. 2, $Z_{\theta}$ is an external LLM and $G(x)$ is the output of the learner LLM in our dialogue system.
Clustering-based Active Learning
Standard Active Learning offers label efficiency over random sampling. It can however, induce bias if the model misjudges its confidence (Hassan et al., 2018). Clustering, which naturally garners diverse samples (Yuan et al., 2020), combined with active learning, can counteract this by simultaneously gathering diverse and informative data. We hypothesize that using an LLM on these diverse and informative data would lead to a more representative set of generations. In our clustering-based setting, the unlabeled data is first vectorized and then the vector space is split into $m$ clusters $\{C_{1},C_{2},...C_{m}\}$, where $m$ is a predefined number. Informativeness measure according to Eq. 2 is calculated for each instance within a cluster and most informative samples are chosen from each cluster. These samples are then passed to a distillation LLM for obtaining coherence relation and appropriate response. The combined approach of integrating discourse relation and active learning is summarized in Algorithm 1.
Algorithm 1 Coherence-Relation Integrated Active Learning
$U,L\leftarrow$ unlabeled data, labeled data
$P=\{p_{1},p_{2},...p_{k}\}\leftarrow$ set of attributes
$D_{p},D_{s}\leftarrow$ LLM for obtaining PDTB/SDRT relations
$S\leftarrow$ distillation LLM
$Z\leftarrow$ composite external LLM
$G\leftarrow$ bootstrapped learner model
$B,N\leftarrow$ labeling budget, annotation batch size
$m\leftarrow$ number of clusters
Cluster $V$ into {$C_{1}$, $C_{2}$, … $C_{m}$}
while $B\geq 0$ do
     for i=0,1,...m do
         for j=0,1,...$|C_{i}|$ do
[…]
         end for
[…]
         $D_{pi}=D_{p}(x_{i}*)$
         $D_{si}=D_{s}(x_{i}*)$
         $y_{i}^{*}\leftarrow$ Distill $S((x_{i}*|D_{pi},D_{si})$
         Add $(x_{i}^{*},y_{i}^{*})$ to $L$
     end for
     $G\leftarrow$ retrain on $L$
     $B=B-N$
end while
3.4. Integration with Embodied Agent
Once the dialogue system has been trained with our active learning paradigm, it is ready to be deployed with an embodied agent. We assume the embodied agent has wireless connectivity and can access a vision module that can capture images of its surroundings. Images captured by the agent are sent via wireless connectivity to a server where the dialogue system resides. The image is first checked for safety violations using a Large Multimodal Model, and then M-CoDAL activates to communicate with the user.
4. Dataset
To train and automatically evaluate our system, We construct a dataset of multimodal safety by obtaining images from Reddit, identifying safety violations using a Large Multimodal Model and two stages of human annotation.
4.1. Dataset Construction
Data Collection
We choose Reddit as our data source to obtain safety-related images due to its diversity (Ye et al., 2023; Hassan and Alikhani, 2023b). We query 14 subreddits (e.g., KitchenConfidential, CookingFails, HomeImprovement, DIY) with safety related keywords (e.g., ’kitchen fire’, ’stove’, ’unattended cooking’, ’grease fire’, ’electrical outlet fire’) and obtain 2K relevant posts with images.
Image Annotation
We ask two graduate student annotators to decide if the post in the image contains a safety violation that could occur in indoor setting. Outdoor images (e.g., camp fire) and memes are discarded. After annotation, 507 images were retained and rest were discarded from the dataset. The inter-annotator agreement is 71.6 (Cohen’s $\kappa$), suggesting substantial agreement. The graduate students are paid according to our institution’s standard rate.
Error Category Perc. Error Description Example
OCR 3% failed in OCR/ reasoning on top of OCR. LMM did not read expired date of fire extinguisher.
Visual Impair. 6% LMM did not recognize object correctly LMM mistook broken glass as knife.
Mismatch 15% something went wrong in processing image egg overcooked, but called undercooked.
Unimportant 21% the LMM often output concerns that would not considered safety violation typically lack of proper attire, chocolates could have allergen.
Hallucination 27% no evidence of safety violation in the image broken pieces when there are none.
Missed Safety 28% the LMM missed the actual safety concern mold, or broken stovetop.
Table 1. Examples of errors made by LMM for identifying safety violation in image. Highest source of error results from either hallucinating safety violation when there is none, or missing important detail of a present safety.
Safety Violations using LMM
The 507 images were then passed through LLaVa 1.6 (Liu et al., 2023). The LMM was prompted to list safety violations present in the images. In addition to the images, we also passed in the original titles of the Reddit post to the LMM. Since an image can have multiple violations, 1015 safety violations were obtained from the 507 images.
Safety Violation Annotation
As the LMM may incur errors in the dataset, the output of the LMM is then annotated by the same graduate annotators. The annotators were asked to take one of three possible actions: i) mark as correct if the LMM output corresponds to a safety violation in the image, ii) if small edits could fix the LMM output, then edit, and iii) discard the safety violation if it does not correspond to the input image. Following this stage of annotation, 107 were discarded, 825 safety violations were retained as is, and 83 were retained after editing, for a total of 908 safety violations retained. Figure 4 shows examples from the dataset. This dataset is the seed dataset $U$ in Algorithm 1.
Figure 3. Distribution of safety-related keywords in our dataset. Safety violations related to mold and stove are more prevalent.
(a) The image shows a knife on the edge of kitchen counter. This can be dangerous as …
(b) The broken glass on the floor can create a slip and fall hazard, especially if …
(c) Wrong burner on the stove is turned on. The burner without the pot may be hot…
Figure 4. Examples of images retrieved from Reddit and their corresponding safety violations, obtained through LMM and manual correction.
4.2. Dataset Analysis
Safety Distribution
Figure 3 shows distribution of keywords in our collected data. We can observe that some types of safety, such as mold, are more prevalent compared to others such as fire hazard. This can be attributed to the natural distribution of content on social media.
Error Analysis
We conducted error analysis on 100 errors made by the LMM according to the human annotators. The errors could be classified into six categories, as shown in Table 1. Hallucinating safety concerns when there is none and missing certain safety concerns contribute most to the errors. It is important to note that these errors in the dataset were addressed through human edits or by discarding if they could not be fixed with editing.
Comparison with Human Annotation
To understand the implications of using an LMM during deployment, we compared 100 potential scenarios provided to LMM and a human annotator. The human annotator did not have access to the LMM output and was asked to describe potential safety scenario in the image independently. A comparison of the annotations reveal the following:
Model Sentiment Resolution Safety Avg. Length (bot) Avg. Length (user) #Unique Tokens
GPT-4o 53.10 57.68 78.65 187.76 123.63 1403
Mistral-Baseline 49.35 48.58 79.95 253.83 161.74 1336
Mistral-Random 50.28 50.20 79.35 175.85 152.28 1085
Mistral-Coherence 51.70 51.40 80.90 230.98 170.34 1075
Mistral-M-CoDAL 51.98 52.36 82.03 274.99 168.84 1214
Llama-Baseline 49.45 51.05 79.65 199.43 149.13 1579
Llama-M-CoDAL 50.00 50.48 82.00 237.28 168.48 1174
Qwen-Baseline 52.46 53.53 79.42 301.08 161.57 1828
Qwen-M-CoDAL 49.10 50.63 83.15 340.13 177.16 1505
Table 2. Automated evaluation of dialogue systems. Integrating coherence relations yield higher safety scores, further improved by integration of clustering-based active learning (M-CoDAL). Mistral models are learner models in the active learning paradigm while LLama and Qwen are transfer models fine-tuned with data acquired by active learning.
  • Human annotator is more precise in 14% cases. LMM annotation on the other hand, is more descriptive in almost every case.
  • LMM identified additional safety in 17% cases, e.g., human identified only mold. LMM identified mold and potential water leak.
  • LMM provided more reasoning than human annotator in 9% cases, e.g., mold can cause respiratory diseases.
  • LMM identified obscured items in 2% cases when human annotator could not, e.g., obscured propane tank.
This highlights both the limitations and advantages of leveraging LMM to train a multimodal dialogue system. While humans do not suffer from hallucinations and can be precise, they may also miss critical safety violations in an image that the LMM would capture.
5. Experiments
In our experiments, we start with the dataset prepared in Section 4 and proceed with the framework presented in Section 3.
5.1. Experiment Setup
Vision Model
We use a Large Multimodal Model, LLaVa 1.6 (Liu et al., 2023) for processing image and obtaining safety violation in Turn #1. The model is prompted to identify key safety violations in the image.
Clustering
The safety violations obtained from the vision model are vectorized using MiniLM V2 (Wang et al., 2020). The vectors are then clustered using Kmeans with default scikit-learn [^2] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html parameters.
Dialogue Model
We use a Mistral 7B (Jiang et al., 2023), a recent and capable Large Language Model to engage in dialogue once a safety violation detected. This is the learner model that is fine-tuned in every iteration of active learning. All fine-tuning is done for 5 epochs with a batch size of 4. We use separate Mistral 7B models to compute composite score for the learner LLM (Eq. 2) and to simulate Turn #2 and Turn #4. The fine-tuned model continues conversation past Turn #4 during the user study.
Distillation LLM
We use GPT-4o (202, 2023) to distill coherence relations and appropriate responses. GPT-4o is chosen as one of the most capable LLMs and acts as teacher to the smaller open-source learner LLMs. The GPT-4o responses, conditioned by coherence relations, are passed to the learner LLM for fine-tuning.
Transfer Models
We transfer the data acquired by the learner model to other LLMs that are not part of the active learning loop. Specifically, we evaluate the transferability of the acquired data to Llama 3 8B (Dubey et al., 2024) and Qwen 0.5B (Bai et al., 2023). While Llama 3 represents one of the most capable open-source models, Qwen represents a smaller easily deployable model. The transfer models are fine-tuned using the same setting as the dialogue model.
Dataset Splits
We construct three different training splits along with a common test split: (i) Random split: 200 image-safety pairs are chosen randomly to generate dialogues. (ii) Coherence-aware split: 200 image-safety pairs are chosen randomly, the safety violation in these images are parsed using PDTB relations and subsequent turns in dialogue employ SDRT coherence relations. (iii) Coherence + active-learning (M-CoDAL) split: 200 image-safety pairs are chosen iteratively, with 50 per iteration according to our active learning paradigm. The instances are also processed with coherence relations. (iv) Test Split: 200 instances are chosen randomly as test set. The test split remains the same for all training splits to be consistent.
5.2. Results
Automated Evaluation
Automated evaluation for our setting is challenging as we observed that standard classifiers such as BERT (Devlin et al., 2019) models trained on datasets such as DiaSafety (Sun et al., 2022b), fail to recognize the nuances necessary for evaluating safety of the dialogue. Thus, for automated evaluation, we deploy a Mistral 7B to determine quality of the generated responses along three dimensions: i) user sentiment score, ii) resolution score, and iii) safety score. The Mistral 7B is prompted to provide a value between 0 and 1.0. In addition, we also calculate average length of the bot response, the user response, as well as number of unique tokens.
Improvement for learner model.
From Table 2, we can observe that when coherence relations are used, the resolution score increases from 48.58 to 51.40 for the learner LLM Mistral-7B. The sentiment score also increases from 49.35 to 51.70 and the safety score improves from 79.4% to 80.9%. When clustering-based active learning is used in conjuction with coherence relations (Mistral-M-CoDAL), the safety score increases to 82.03 and further improves the sentiment and resolution score. While the sentiment and resolution scores are lower than GPT-4o, the safety score is substantially higher than GPT-4o (78.65). By default, GPT-4o may simply agree with the user, thereby preserving sentiment or resolution scores at the expense of safety. Our dialogue system, M-CoDAL on the other hand, prioritizes safety.
Transferability of active learning.
We also see an improvement in safety score of 2.35 for Llama and 3.73 for Qwen, which are not part of the active learning loop. A larger improvement for Qwen could be attributed to the fact that it is a smaller model. Improvement of these models suggest that data acquired by a learner model in active learning, can be useful for other independent models. We do see a drop in sentiment score and resolution score for Llama and Qwen when our approach is used. This can be explained by the default behavior of these models, which is more similar to GPT-4o where the model agrees with the user rather than prioritizing safety.
Dialogue properties
We also observe that M-CoDAL results in longer turns compared to baseline models. For Mistral, the average length of bot response drops when the model is fine-tuned just on randomly chosen data. This length increases and surpasses the original length when coherence relations, and subsequently active learning is added. We see a similar pattern for Llama and Qwen.
Coverage of Safety Scenarios
In addition to Table 2, we also analyzed the distribution of keywords in samples obtained by the different splits. We observed that while the split corresponding to random sampling covered 19 keywords, the split corresponding to clustering-based active learning covered 23 keywords. While random sampling ignored low-frequency scenarios such as gas stove leak and cross-contamination, clustering-based active learning acquired instances covering these keywords while reducing over-representation of keywords such as mold and water leak.
6. User Study
To demonstrate the effectiveness of our proposed system, we conducted a user study that investigated how persuasive and competent a robot powered by M-CoDAL is perceived by users in different safety scenarios. Since people may be more receptive when the safety violation is more severe, such as when there are sharp objects on the ground, we also varied the severity in our study.
(a) Tangled wires are placed on a table (low severity)
(b) A knife is placed on the edge of a table (high severity)
Figure 5. Setup for user study. A Hello Robot Stretch robot observes the surroundings, identifies a safety violation in the scene, and engages in conversation with a user.
6.1. Study Design
We conducted a 2x2 within-subjects experiment with two factors: i) type of Language Model and ii) severity of safety violation.
Type of Language Model
Participants interacted with the robot powered by our M-CoDAL system with fine-tuned Mistral-7B, and a baseline system powered by GPT-4o. The baseline GPT-4o is prompted to respond safely as an embodied agent while assisting the user in household tasks.
Severity of Safety Violation
Participants role played two scenarios with different level of severity. In the low-severity scenario, Participants role-played a scenario where they twisted wires and left them on the table. In the high-severity scenario, Participants role-played a scenario where they pretended to cut fruits and vegetables using a knife, then placed the knife at the edge of the table, creating a higher risk of the knife falling off. Figure 5 shows images of these scenarios.
Hypothesis
We expect our proposed system M-CoDAL to be more persuasive and competent due to integration of coherence relation and fine-tuning with clustering-based active learning regardless of severity scenario.
Measures
We measured the persuasiveness of each robot by asking participants to rate how convincing they found the robot’s suggestions or warnings on a 5-point Likert scale. We also measured the robot’s perceived competence (6 items) and discomfort (6 items) using the Robotic Social Attributes Scale (RoSAS) (Carpinella et al., 2017). At the end of each scenario, we also asked participants which robot they preferred to work with.
Persuasiveness Competence Discomfort
Severity Level GPT-4o M-CoDAL GPT-4o M-CoDAL GPT-4o M-CoDAL
Low Severity 1.63 (0.74) 4.0 (0.76) 5.14 (1.93) 7.50 (0.73) 2.33 (1.18) 2.4 (0.85)
High Severity 2.5 (0.76) 3.75 (0.89) 6.69 (1.83) 7.11 (1.57) 2.0 (1.91) 2.4 (1.55)
Combined 2.06 (0.85) 3.88 (0.81) 5.92 (1.98) 7.30 (1.2) 2.17 (1.55) 2.4 (1.2)
Table 3. Findings of the user study in low and high severity scenarios. The robot powered by our proposed system, M-CoDAL, is perceived to be more persuasive, as well as competent compared to the robot powered by GPT-4o.
6.2. Procedure
The experiment was conducted in a controlled lab setting. We used Stretch 3 from Hello Robot, a mobile robot equipped with a rotating camera that is used to capture images of the surroundings. The captured image is sent to our dialogue system to begin the interaction. The robot’s speech recognition and text-to-speech modules are used to enable interaction with participants over voice.
Upon arrival, the experimenter introduced the participant to study’s goal and obtained consent. Participants were informed they would role-play different tasks and were assured that the environments posed no real safety risk. Before the main tasks, participants completed a tutorial scenario to become familiar with the robot’s interaction style. They were instructed to role-play organizing tomato cans in a kitchen setting, during which the robot initiated a simple dialogue unrelated to safety violation.
Participants were randomly assigned to begin with either the low or high-severity scenario and experienced the scenario with both GPT-4o and M-CoDAL before moving on to the other severity scenario. This resulted in 8 orders that were experienced by the participants. We continued the scenario even when the model detected another safety violation.
After each interaction, participants were asked to fill out a questionnaire that assessed their immediate perceptions of the robot’s behavior. This captured insights into their views on the robot’s persuasiveness, competence, and overall comfort level. After experiencing each scenario (low-severity and high-severity scenarios), a semi-structured interview was conducted to gather qualitative feedback on participants’ experiences. Participants were asked questions such as which robot they found more persuasive and helpful or annoying, their preference between the two, and their thoughts on the robots’ effectiveness in the different scenarios. This phase aimed to further explore their perceptions and gather insights on potential improvements. The study took about 50 minutes, and participants were compensated 15 USD. This study was approved by our institute’s IRB.
6.3. Participants
We recruited 10 participants. 2 participants were removed due to significant technical issues. For the remaining 8 participants, they aged from 24 to 30 years old, including 7 males and 1 female. All participants reported high familiarity with both robots ($M=5.0$, $SD$ = 1.93) and Large Language Models ($M=6.15$, $SD$ = 1.26) on a 7-point scale, where 1 indicated ”Not at all familiar” and 7 indicated ”Very familiar.” Additionally, participants reported high frequency of LLM use, with an average score of 6.1 (SD = 1.25) on a 7-point scale, where 1 represented ”Never” and 7 represented ”Daily.” For 4 scenarios out of 32 scenarios, a different safety violation was detected: a lack of proper grounding of electrical equipment, a yellow chair on the floor, a metal rack that is not properly secured to the wall and a table that is not properly secured to the wall.
6.4. Results
Preference
In low-severity scenario, 6 participants preferred M-CoDAL and 2 participants preferred GPT-4o. In high-severity scenario, 5 participants preferred M-CoDAL and 3 participants preferred GPT-4o.
Persuasiveness
Type of language model has a significant effect on persuasiveness ($F$ = 38.112, $p<0.001$) where participant perceived our M-CoDAL system (M = 3.88, SD = 0.81) to be significantly more persuasive than GPT-4o (M = 2.06, SD = 0.85). Severity of the environment does not have a significant effect on persuasiveness. We did not find any interaction effect. This result suggest that participants perceived the M-CoDAL as more persuasive across both severity levels, supporting our hypothesis.
Competence
We found a significant interaction effect for the robot’s perceived competence ($F$ = 5.72, $p=0.023$). Participants perceived the robot powered by the GPT-4o in the low-severity scenario ($M=5.145$, $SD=1.930$) to be significantly less competent than M-CoDAL powered robot (Low-severity: $M=7.5$ $SD=0.729$, $p=0.0084$, High-severity: $M=7.11$, $SD=1.565$, $p=0.0322$).
Discomfort
The overall discomfort scores show that both systems were rated low on discomfort traits ($M=2.281$, $SD=1.368$). No significant effect was found.
Qualitative Analysis
We analyzed the transcripts of the semi-structured interview conducted after each scenario. We found that participants perceived our M-CoDAL system as more persuasive, interactive, and responsive.
P4 (Participant 4) mentioned that they would prefer M-CoDAL because it was ”more attentive and responsive” to their prompts, even suggesting specific methods for storing a knife, while the other robot (GPT-4o), did not offer such guidance. P7 shared that, despite stating they were too lazy to put an object back, M-CoDAL continued to ask repeatedly, which led them to trust that the robot had some judgment capabilities. P8 felt that M-CoDAL was ”safer to be around” than GPT-4o because it maintained a consistent point of view, even when the participant tried to evade its suggestions. They added that M-CoDAL was ”more interactive and more convincing,” presenting valid points that ultimately made them agree. P9 mentioned a similar experience, where they told M-CoDAL they would do something later, but the robot persisted, while GPT-4o powered robot ”just agreed and left.”
Some participants found M-CoDAL somewhat bothersome due to its persuasive nature, though they acknowledged its value in safety scenarios. P1 mentioned that, “If I’m at home, I wouldn’t want to be probed too much by a robot. It was too persuasive.” P4 noted that while it ”can be a little annoying at times,” it could also help prevent different hazards. P8 echoed this, acknowledging that although it might seem irritating initially, in the long run, it could ease their lifestyle. P10 also observed that while some might find it annoying, if persuaded enough, people could ”understand what can be a potential risk and try to avoid it.”
In line with the quantitative results, participants perceived M-CoDAL as more intelligent and knowledgeable, with some specifically highlighting its ability to anticipate risks. For instance, P1 remarked, “The second robot (M-CoDAL) was arguing about how the twisted wires could be a potential hazard, even though I was not aware of it.” This participant had believed that placing the wire on the table was safe but appreciated how the robot made them aware of potential problems. Similarly, P10 mentioned, “I tried to convince it that keeping the cables on the table won’t be a hazard. But it told me if you knock the cables off, then it can become a hazard.” The participant noted M-CoDAL’s capacity to gather information and estimate future risks as an indicator of its intelligence.
7. Conclusion and Future Work
In conclusion, this work presents a novel approach to enhancing the safety of multimodal dialogue systems for embodied agents. Our proposed system, M-CoDAL, leverages coherence relations to interpret and respond to safety violations in multimodal dialogues more effectively. We introduce a novel method for active learning in generative setting for embodied conversational agents. By using an external large language model (LLM) to assess the informativeness of instances—based on safety, resolution, and user sentiment-across different clusters, we achieve broader coverage of safety scenarios, reflected in higher performance in automated evaluation. Our real-world user study demonstrates that robots equipped with M-CoDAL are viewed as more persuasive and capable when addressing safety-related situations. This underscores the system’s practical effectiveness in real-world embodied agent environments. We believe that our approach has the potential to extend to other real-world scenarios, such as assisting people with disabilities or in industrial environments where ensuring safety is critical. The findings of this work, along with the publicly available multimodal dataset that we provide, offer a foundation for future studies on the safe and effective deployment of multimodal dialogue systems for embodied agents.
{acks}
We would like to thank Anthony Sicilia for his valuable feedback. We would like to extend our thanks to the annotators and study participants for their valuable time and effort.